suppose you flip a oin 3 times what is the probability that at least once the coin comes up head
Answer by stanbon(75887)

suppose you flip a oin 3 times what is the probability that at least once the coin comes up head
---------------------------
P(at least one head) + P(no heads) = 1
P(at least one head) = 1 - P(no heads)= 1 - (1/2)^3 = 1-(1/8) = 7/8
=============================================
Cheers,
Stan H.